# Associative Transformer

### v3

@@ Appendix E Efficacy of the Bottleneck Attention Balance Loss
The Bottleneck Attention Balance Loss facilitates the learning of priors that attend to diverse sets of patches. We demonstrate the efficacy by visualizing the bottleneck attention scores (Figure 7) and the selected patches by the bottleneck attention (Figure 8). We employed as a metric for patch diversity the ratio of distinct patches in all the selected patches.
@@
Figure 7: The Bottleneck Attention Balance Loss facilitates the selection of more diverse patches from various input positions.
@@
Figure 8: Selected patches by the bottleneck attention in CIFAR-10, demonstrating an apparent increase in the diversity of selected patches as training progresses.
Appendix F Hopfield Networks Energy
@@
Moreover, the information retrieval is based on a continuous Hopfield network, where an input state converges to a fixed attractor point within the associative memory of the Hopfield network. Usually, any input state that enters an attractor’s basin of attraction will converge to that attractor. The convergence results in a decreased state energy with respect to the stored attractors in the memory. All patches reach their minimum at the same time and the global energy in Equation 7 is guaranteed to decrease. To quantitatively measure the amount of energy reduction during the information retrieval process in the Hopfield network, we computed an input state’s energy before and after it was reconstructed. A successful retrieval results in substantial reduction in the state energy (Figure 9).

### v4

@@ Appendix G Efficacy of the bottleneck attention balance loss
The Bottleneck Attention Balance Loss facilitates the learning of priors that attend to diverse sets of tokens. We demonstrate the efficacy by visualizing the bottleneck attention scores (Figure 6) and the corresponding patches of the selected tokens by the bottleneck attention (Figure 7). We employed as a metric for patch diversity the ratio of distinct tokens in all the selected tokens.
@@
Figure 6: The Bottleneck Attention Balance Loss facilitates the selection of more diverse tokens from various input positions.
@@
Figure 7: The corresponding patches of the selected tokens by the bottleneck attention in CIFAR-10, demonstrating an apparent increase in the diversity of selected tokens as training progresses.
We visualized the attention scores over different tokens in Figure 8. Despite employing tokens across different batch samples, each prior eventually learned to extract from similar patch input positions, demonstrating emergent spatial specialization.
Figure 8: Prior-guided attention visualization highlights selected tokens, with each prior having learned to extract from similar patch input positions across different samples. These priors exhibit emergent specialization, focusing on specific spatial areas in an image (using the first global workspace layer of AiT-Small in CIFAR-10).
Appendix H Hopfield networks energy
@@
Moreover, the information retrieval is based on a continuous Hopfield network, where an input state converges to a fixed attractor point within the associative memory of the Hopfield network. Usually, any input state that enters an attractor’s basin of attraction will converge to that attractor. The convergence results in a decreased state energy with respect to the stored attractors in the memory. All tokens reach their minimum at the same time and the global energy in Equation 5 is guaranteed to decrease. To quantitatively measure the amount of energy reduction during the information retrieval process in the Hopfield network, we computed an input state’s energy before and after it was reconstructed. A successful retrieval results in substantial reduction in the state energy (Figure 9).